OpenStudy (lgbasallote):

Prove that if n = ab, where a and b are positive integers, then \(a \le \sqrt n\) or \(b \le \sqrt n\)

OpenStudy (chihiroasleaf):

prove by contrapositive

OpenStudy (lgbasallote):

hmm....and that means...?

OpenStudy (anonymous):

The contrapositive would be: If \(a > \sqrt{n} and b > \sqrt{n}\) then \(n \ne ab\).

OpenStudy (chihiroasleaf):

show that \[a >\sqrt{n}\] and \[b>\sqrt{n}\] implies \[n \neq ab\]

OpenStudy (lgbasallote):

hmm 3 posts....

OpenStudy (anonymous):

Shhh, there's no edit button.

OpenStudy (anonymous):

And for some reason, no preview. When did that happen?

OpenStudy (lgbasallote):

okay... so... \[a^2 > n\] \[b^2 > n\] then.... \[a^2 - b^2 = 0\]

OpenStudy (lgbasallote):

what am i doing with my life...

OpenStudy (anonymous):

How do you get \(a^2-b^2 = 0\)? That's not right.

OpenStudy (lgbasallote):

\[a^2 - b^2 > 0\] you know what i mean\t

OpenStudy (anonymous):

Still not right.

OpenStudy (lgbasallote):

??????!!!!!!!!!!

OpenStudy (chihiroasleaf):

assume that \[a>\sqrt{n}\] and \[b>\sqrt{n}\] what do you get for ab ?

OpenStudy (anonymous):

b could be larger than a. That is not restricted by your assumption, nor does it need to be.

OpenStudy (lgbasallote):

fine. \[a > \sqrt n\] \[b > \sqrt n\] \[a - b > 0\] happy?

OpenStudy (anonymous):

Not super happy. I'm not sure why you care so much about the difference of a and b, when the proof is mostly concerned with their product.

OpenStudy (lgbasallote):

.........

OpenStudy (chihiroasleaf):

your goal is to show that \[n \neq ab\] why do you subtract a-b?

OpenStudy (lgbasallote):

because i can...

OpenStudy (anonymous):

lol okay.

OpenStudy (lgbasallote):

.....i hate math.....

OpenStudy (chihiroasleaf):

now try to multiply a and b, what do you get? :D

OpenStudy (lgbasallote):

if a > sqrt n and b > sqrt n.... then \[ab > \sqrt n \cdot \sqrt n\] \[ab > n\]

OpenStudy (lgbasallote):

is that allowed?

OpenStudy (anonymous):

That's the basic logic of it.

OpenStudy (chihiroasleaf):

yes..., :D

OpenStudy (anonymous):

You start with \(a > \sqrt{n}\) Multiply both sides of the equation by b. Do we have to change the direction of the inequality? No, because we are assuming that b is positive. \(ab > \sqrt{n}b\)

OpenStudy (anonymous):

Now, because \(b > \sqrt{n}\) \(\sqrt{n}b > \sqrt{n}\sqrt{n} = n\)

OpenStudy (anonymous):

Therefore, ab > n.

OpenStudy (lgbasallote):

that's really allowed?

OpenStudy (anonymous):

Kind of. You have to do those middle steps, and it's important to understand that multiplying both sides by b does not change the direction of the inequality only because we are assuming b is positive.

OpenStudy (anonymous):

Allow me to demonstrate. 2 > 1 Shall I multiply both sides by x? Sure, let's do it. 2x > x Seems true, right? Well, it's true if x is positive. x=5 10 > 5 What about if x is negative? x=-5 -10 > -5 Aw sadness. Not true anymore.